OpenStudy (anonymous):

The surface area of a soup can represents the total area of the can's exterior and may be found by the formula S.A. = 2πrh + 2πr2. A soup company designing its own cans wants to determine the height of its cans with a given surface area and radius. Help the company by solving the surface area formula for h.

OpenStudy (anonymous):

Probably the first step is SA = 2 pi r ( h + r )

OpenStudy (anonymous):

my answers are h= 2 pi r/SA+2 pi r h= 2 pi r/SA-2 pi r h= SA+2 pi r/2 pi r h=SA-2 p r/2 pi r

OpenStudy (anonymous):

*answer choices

OpenStudy (anonymous):

I made a little mistake above. S.A. = 2πrh + 2πr^2 2πrh = SA - 2πr^2 h = SA/2πr + r None of the answers looks correct.

OpenStudy (anonymous):

these are the answers i can choose from